OpenStudy (anonymous):

Solve for y: 2/3 y + 2 = 8 y = 11 y = -11 y = -9 y = 9

OpenStudy (anonymous):

y-9

OpenStudy (anonymous):

y-9

OpenStudy (anonymous):

take 2 away from each side 2/3y = 8-2 multiply both sides by 3/2 y = 3/2*(6) y= 3*3 y=9

OpenStudy (anonymous):

nicole can you clarify the formula please